Hello.
Could someone please help me complete the table for the function.
f(x)=x^2
x= 0, + or - 1, + or - 2, + or - 3.
y=
thanks.

Answer by Alan3354(69443)

Could someone please help me complete the table for the function.
f(x)=x^2
x= 0, + or - 1, + or - 2, + or - 3.
y=
y = f(x)
==================
f(0) = 0^2 = 0
f(-1) = (-1)^2 = 1
f(20) = 20^2 = 400
See how that works?